OpenStudy (anonymous):

Is the simplified form of 2square root of 3 • square root of 12 rational? Yes No

OpenStudy (timaashorty):

\[2 \sqrt 3 * \sqrt 12 = \] \[2\sqrt3 * \sqrt 4 * \sqrt 3 = \] \[2 \sqrt 3 * 2 \sqrt 3 =\] \[4 * 3\] Therefore, Yes it's rational.

OpenStudy (anonymous):

Thank you so much!! I greatly appreciate you helping me on this.

OpenStudy (timaashorty):

no problem and also, welcome to Openstudy (:

OpenStudy (anonymous):

Thank you. Haha. Can you help me with a few more? I'm not really understanding these

OpenStudy (timaashorty):

I can try

OpenStudy (anonymous):

4 square root of 2 end root plus 7 square root of 2 end root minus 3 square root of 2 a.)2 square root of 8 b.)8 square root of 2 c.)8 square root of 6 d.)6 square root of 8

OpenStudy (timaashorty):

\[4 \sqrt 2 + 7 \sqrt 2 - 3 \sqrt 2 = 8\sqrt 2\] B. 8 square root of 2

OpenStudy (anonymous):

Ohhh. I did it wrong then. I got a... Haha. Ooppsiee. ALrighty heres the next one 5 square root of 7 end root plus 12 square root of 6 end root minus 10 square root of 7 end root minus 5 square root of 6 5 square root of 14 end root minus 7 square root of 12 5 square root of 7 end root minus 7 square root of 6 7 square root of 12 end root minus 5 square root of 14 7 square root of 6 end root minus 5 square root of 7

OpenStudy (timaashorty):

\[5 \sqrt 7 + 12 \sqrt 6 - 10 \sqrt 7 - 5 \sqrt 6 = 7 \sqrt 6 - 5 \sqrt7\] D 7 square root of 6 end root minus 5 square root of 7

OpenStudy (timaashorty):

Can you post your other question in a new post?

OpenStudy (anonymous):

Sure! :)

OpenStudy (timaashorty):

thnks [: